# A new approach to principal–agent problems with volatility control

### v1

@@ 1 Introduction
In terms of theoretical developments, the most significant breakthrough is thanks to Cvitanić, Possamaï, and Touzi (2018) [9], who have developed a general and rigorous theory that allows to address a wide spectrum of principal–agent problems, especially when the agent can also control the volatility of the output process. As explained in [9], moral hazard problems in continuous-time in which the agent controls the volatility of the output are notoriously harder to study. Indeed, when the agent is allowed to also control the volatility of the output process, his continuation utility cannot be related to the solution of a BSDE anymore, one has to rely now on the theory of backward stochastic differential equations of the second order (2BSDE for short), introduced by Soner, Touzi, and Zhang (2012) [30] and further developed by Possamaï, Tan, and Zhou (2018) [28]. One important observation made in [9] which will also be at the heart of our approach is that, if the principal has access to the trajectory of the output process $X$, then she can deduce its quadratic variation process $\langle X\rangle$ in a pathwise manner, as demonstrated by Bichteler (1981) [4]. This motivates the authors to introduce a new form of contracts, namely
@@
Such contracts are indexed on $X$ as before, but also on its quadratic variation $\langle X\rangle$ through an additional parameter $\Gamma$, which should incentivise the agent to perform appropriate efforts on the volatility. Note that the agent’s Hamiltonian ${\cal H}$ here should be different from the one introduced in the previous contract form, precisely to take into account volatility control. When restricting the study to this form of contracts, it is straightforward to show, mostly using Ito’s formula, that the agent’s optimal effort will naturally be given by the maximiser of the Hamiltonian. One can then proceed by solving the principal’s optimal control problem, namely by writing the Hamilton-Jacobi-Bellman (HJB for short) equation corresponding to her stochastic control problem, which becomes relatively standard under the restriction to the previous form of contracts. These first two steps are therefore quite straightforward using standard tools from stochastic control theory. However, the third and last step requires to prove that the restriction to such form of contracts is without loss of generality. The approach provided in [9] to prove this main result relies on the recent and sophisticated theory of 2BSDEs. The aim of this paper is precisely to present an alternative approach to solving this type of problems, which avoids relying on this theory. For the sake of comparison, we will study the same general model, whose original formulation will be recalled in Section 2.
@@
Despite the challenges raised when studying principal–agent problems with volatility control, especially as the optimality of revealing contracts relies on deep results from the recent theory of 2BSDEs, it has subsequently been extended and applied in many different situations. We can mention in a non-exhaustive way the applications related to finance by the same authors in [8], by Cvitanić and Xing (2018) [7] or more recently by Baldacci and Possamaï (2022) [3]; the works by Aïd, Possamaï, and Touzi (2022) [1], Élie, Hubert, Mastrolia, and Possamaï (2021) [11] and Aïd, Kemper, and Touzi (2023) [2] for applications related to the energy sector; as well as other various extensions, e.g., the works of Mastrolia and Possamaï (2018) [22], Hernández Santibáñez and Mastrolia (2019) [14], Hu, Ren, and Touzi (2019) [17], Hubert (2023) [18] or Keppo, Touzi, and Zuo (2024) [21]. Naturally, the study of applications fitting into the general framework proposed in [9] can be done in a fairly straightforward way, directly using the results, as done in [8, 1]. Nevertheless, the recent research mentioned above has also attempted to generalise principal–agent problems, especially to consider many agents (see [18, 2]), or even a continuum (see [11]). If one follows the approach developed in [9], these extensions require to introduce the so-called multidimensional or mean-field 2BSDEs, and study, or assume, the well-posedness of such systems. However, while the theory on (one-dimensional) 2BSDEs is sufficiently developed to address relatively general principal–agent problems, there is currently no definition nor existence results on multidimensional/mean-field 2BSDEs. This further motivates us to present, in this paper, an alternative approach to solving the principal–agent problem studied in [9]. We believe that this novel approach, which avoids relying on the theory of 2BSDEs, will greatly facilitate the study of relevant generalised principal–agent problems, especially with many agents, or non-continuous output processes.
@@
Before introducing the approach developed in this paper, we wish to provide more intuition on the problem under consideration via a short detour through the so-called first-best case. When studying principal–agent problems with moral hazard, it is quite common to compare the results with this benchmark scenario, in which the principal can directly optimise the agent’s effort. This problem is usually easier to solve, as one only has to consider the principal’s stochastic control problem under the agent’s participation constraint. It may happen in the moral-hazard case, sometimes also referred to as second-best, that the principal can actually deduce, through her observation, the effective efforts of the agent. Under this assumption, the two problems are then equivalent if the principal can design a contract which strongly penalises the agent whenever he deviates from the optimal effort recommended by the principal. Typically, the principal will offer to the agent a minimal [^1] Minimal in the sense that it would just be sufficient to ensure the participation of agent. compensation if he follows the recommendation, and threaten him with a high penalty if he deviates from the optimal effort. This can usually be done if no strong assumptions are made on the contract, but can fail to be implemented in cases for example with limited liability, in which the contract is required to be positive.
@@
As highlighted before, the approach developed by Cvitanić, Possamaï, and Touzi (2018) [9] relies in particular on the observation that, if the principal have access to the trajectory of the output process $X$, she can actually compute its quadratic variation $\langle X\rangle$ pathwise, and thus design contracts of the form (2.11), indexed on $X$ and $\langle X\rangle$ through sensitivity parameters $Z$ and $\Gamma$ to be optimised. The main idea behind our reformulation of the original principal–agent problem with volatility control is that, since the principal can deduce the quadratic variation $\langle X\rangle$, she could probably force the agent to choose his efforts in order to achieve an optimal quadratic variation. In other words, the idea here is to study an alternative ‘first-best’ formulation of the original problem, in which the principal enforces a contract $\xi$ as well as a quadratic variation process. The agent should then optimise his effort in order to maximise his expected utility given such contract $\xi$ as before, but now the admissible set of efforts is restricted to those achieving the quadratic variation imposed by the principal. This ‘first-best’ alternative formulation of the problem is described with more details in Section 3.1. When carefully studying the agent’s problem in this reformulation of the original problem, one can remark that, since the quadratic variation is a priori fixed by the principal, the agent’s weak formulation boils down to choosing equivalent probability measures by means of Girsanov theorem. For a fixed quadratic variation process and any contract $\xi$, the agent’s continuation utility can therefore be represented by means of a BSDE, instead of a 2BSDE in the original formulation. More informally, in this reformulated problem, one can apply the so-called ‘Sannikov’s trick’ to obtain a relevant representation of the contract $\xi$, which is then similar to the optimal form of contracts in principal–agent problems with drift control only, namely (1.1). Finally, the principal’s problem boils down to a more standard stochastic control problem, in the sense that she should maximise her expected utility by optimally choosing the contract’s parameters, namely a constant $y_{0}\in\mathbb{R}$ and an appropriate process $Z$, as well as the quadratic variation process.
We wish to insist on the fact that the aforementioned approach to solve the ‘first-best’ alternative formulation of the original problem is quite straightforward, as it follows the ‘Sannikov’s trick’. More importantly, we establish that the principal’s value in this reformulated problem coincides with the one obtained in [9]. More precisely, from the principal’s point of view, optimising over $\Gamma$ in the contract form (1.2) or directly optimising over admissible quadratic variation lead to the exact same results. In particular, the agent’s optimal effort as well as his utility remain unchanged in both formulations. This result is actually quite intuitive: as contracts of the form (1.2) introduced in [9] are optimal when the principal observes the quadratic variation, they should allow her to achieve the expected utility she would have if she could directly control the quadratic variation process. In particular, this result confirms that contracts of the form (1.2) are also optimal when the agent’s efforts on the volatility can actually be determined by the principal through her observation of the quadratic variation. This equivalence between the two formulations is detailed in Section 3.3, using penalisation contracts.
@@
As already mentioned before, the moral-hazard case is equivalent to its first-best counterpart if the principal can deduce the agent’s control from her observations and design a penalisation contract. Indeed, even if the principal observes the agent’s control, she may not have enough bargaining power to actually force the agent to perform the optimal controls. Showing existence of penalisation contracts thus ensures that the two frameworks are equivalent. Similarly, in our problem, even if the principal observes the quadratic variation of the output process, she may not be able to actually force the agent to follow the recommended quadratic variation. We therefore need to show the existence of such penalisation contracts, which allow the principal to achieve her ‘first-best’ utility even if she cannot directly control the quadratic variation. It is quite clear from the previous reasoning that the contract form (1.2) introduced in [9] will perfectly serve this purpose. We nevertheless wish to emphasise that, although we will use the contract form (1.2) as penalisation contracts, we never, with our approach, use the 2BSDE theory to prove optimality of such contracts, as it naturally comes from the fact that these contracts achieve the best possible utility for the principal. In other words, instead of proving via the theory of 2BSDE that considering contracts of the form (1.2) is without loss of generality from the principal’s point of view, one can simply show that these contracts allow her to reach her ‘first-best’ value.
@@
To summarise, this paper provides an alternative approach to solve principal–agent problems with volatility control. More precisely, we first solve the ‘first-best’ formulation of the original problem by means of the theory of BSDEs, and then show that the contract form studied in [9] allows to achieve this first-best scenario, and is therefore naturally optimal. In other words, this alternative approach allows to prove that the contract form introduced in [9] is optimal, relying only on the theory of BSDEs. Although this approach ‘only’ achieves the same results as those presented by Cvitanić, Possamaï, and Touzi (2018) [9], we believe that this alternative, simpler approach will enable better dissemination of continuous-time principal–agent problems with volatility control in various fields across mathematical finance, economics, management sciences, operation research; fostering their use in many relevant applications. Furthermore, as this new approach only relies on the theory of BSDEs, we believe that it will facilitate the study of relevant extensions of principal–agent problems, for example when considering such a Stackelberg game with many agents, as well as its mean-field limit, when the principal can contract with a continuum of agents. This more straightforward approach could also simplify the extension of principal–agent problems to more general non-continuous output processes.
@@ 2 General model
In order to highlight that our new approach allows to obtain the same results as in [9], we focus here on the exact same model. We believe that this model is sufficiently general to convince the reader of the relevance of our novel approach, and that it could then be extended in a (relatively) straightforward way to solve more complex principal–agent problems.
@@ 3. $(iii)$
as well as for any $\nu\in\widebar{\cal U}$, the set of corresponding probability measures $\widebar{\cal P}(\nu):=\{\mathbb{P}\in\widebar{\cal P}:\nu\in\widebar{\cal U}(\mathbb{P})\}$. To avoid technical considerations (see Remark 2.2 below), we further restrict the admissible controls to those that induce a unique weak solution to (2.3), i.e.:
@@
The underlying assumption in (2.4) that each admissible effort $\nu\in{\cal U}$ induces a unique weak solution to SDE (2.3) is not enforced in [9] but is considered here to avoid technical considerations regarding the martingale representation property. We refer to Remark 3.3 for further discussions on this point, but wish to justify here that this assumption is not far-fetched. Indeed, it is relatively common in the study of principal–agent problems to at least assume that the martingale representation is valid (see Hernández Santibáñez (2022) [13, Assumption 2.2. and Remark 2.3] for example), and weak uniqueness for SDE (2.3) is a sufficient and more explicit condition for such representation to hold. Furthermore, as one of the main objectives of principal–agent problems is to determine ‘revealing’ contracts, i.e. contracts that reveal the agent’s optimal efforts, it is actually questionable to consider efforts that would induce several weak solutions for SDE (2.3), especially in view of applications.
@@ 2.2 The original principal–agent problem
$V_{\rm A}(\xi):=\sup_{\mathbb{P}\in{\cal P}}J_{\rm A}(\xi,\mathbb{P}),\;\text{ with }\;J_{\rm A}(\xi,\mathbb{P}):=\mathbb{E}^{\mathbb{P}}\bigg{[}{\cal K}_{\rm A}^{\mathbb{P}}(T)\,\xi-\int_{0}^{T}{\cal K}_{\rm A}^{\mathbb{P}}(s)c\big{(}s,X,\nu_{s}^{\mathbb{P}}\big{)}\mathrm{d}s\bigg{]},\quad\mathbb{P}\in{\cal P},$ (2.5)
@@ 1. $(i)$
    the measurable cost function $c:[0,T]\times{\cal C}([0,T],\mathbb{R})\times U\longrightarrow\mathbb{R}$ is such that for any $u\in U$, the map $(t,x)\longmapsto c(t,x,u)$ is $\mathbb{F}$-optional, and satisfies the following integrability condition:
    $\sup_{\mathbb{P}\in{\cal P}}\mathbb{E}^{\mathbb{P}}\bigg{[}\int_{0}^{T}\big{|}c\big{(}t,X,\nu^{\mathbb{P}}_{t}\big{)}\big{|}^{p}\mathrm{d}t\bigg{]}<+\infty,\;\text{for some }p>1;$ (2.6)
@@ 2. $(ii)$
    where $k_{\rm A}:[0,T]\times{\cal C}([0,T],\mathbb{R})\times U\longrightarrow\mathbb{R}$ is assumed to be measurable, bounded, and such that for any $u\in U$, the map $(t,x)\longmapsto k_{\rm A}(t,x,u)$ is $\mathbb{F}$-optional.
@@
As mentioned in [9, Remark 2.1], the model and the approach can be extended in a straightforward way to account for risk-aversion in the agent’s utility function. More precisely, in the agent’s expected utility (2.5), $\xi$ can be replaced by $U_{\rm A}(\xi)$ for an appropriate utility function $U_{\rm A}$, modulo the modification of the integrability condition (${\rm I}^{p}_{\rm A}$). As we choose to follow closely the model considered in [9], we do not consider any utility functions here, but we will address the case of CARA—exponential—utility functions through the examples in Section 4.
@@
We continue to follow the framework described in [9] by defining the principal’s optimisation problem as follows:

### v2

@@ 1 Introduction
In terms of theoretical developments, the most significant breakthrough is thanks to Cvitanić, Possamaï, and Touzi (2018) [9], who have developed a general and rigorous theory that allows to address a wide spectrum of principal–agent problems, especially when the agent can also control the volatility of the output process. As explained in [9], moral hazard problems in continuous-time in which the agent controls the volatility are notoriously harder to study. Indeed, in this case, his continuation utility cannot be related to the solution of a BSDE anymore, one has to rely now on the theory of backward stochastic differential equations of the second order (2BSDE for short), introduced by Soner, Touzi, and Zhang (2012) [30] and further developed by Possamaï, Tan, and Zhou (2018) [28]. One important observation made in [9] which will also be at the heart of our approach is that, if the principal has access to the trajectory of the output process $X$, then she can deduce its quadratic variation process $\langle X\rangle$ in a pathwise manner, as demonstrated by Bichteler (1981) [4]. This motivates the authors to introduce a new form of contracts, namely
@@
These contracts are indexed on $X$ as before, but also on its quadratic variation $\langle X\rangle$ through an additional parameter $\Gamma$, which would incentivise the agent to perform appropriate efforts on the volatility. Note that the agent’s Hamiltonian ${\cal H}$ here should be different from the one introduced in the previous contract form, precisely to take into account volatility control. When restricting the study to this form of contracts, it is straightforward to show, mostly using Ito’s formula, that the agent’s optimal effort will naturally be given by the maximiser of the Hamiltonian. One can then proceed by solving the principal’s optimal control problem, namely by writing the Hamilton-Jacobi-Bellman (HJB for short) equation corresponding to her stochastic control problem, which becomes relatively standard under the restriction to the previous form of contracts. These first two steps are therefore quite straightforward using standard tools from stochastic control theory. However, the third and last step requires to prove that the restriction to such form of contracts is without loss of generality. While the approach provided in [9] to prove this main result relies on the theory of 2BSDEs, the aim of this paper is precisely to present an alternative approach which avoids relying on this sophisticated theory. For the sake of comparison, we will study the same general model, whose original formulation will be recalled in Section 2.
@@
Despite the challenges raised when studying principal–agent problems with volatility control, especially as the optimality of revealing contracts relies on deep results from the recent theory of 2BSDEs, it has subsequently been extended and applied in many different situations. We can mention in a non-exhaustive way the applications related to finance by the same authors in [8], by Cvitanić and Xing (2018) [7] or more recently by Baldacci and Possamaï (2022) [3]; the works by Aïd, Possamaï, and Touzi (2022) [1], Élie, Hubert, Mastrolia, and Possamaï (2021) [11] and Aïd, Kemper, and Touzi (2023) [2] for applications related to the energy sector; as well as other various extensions, e.g., the works of Mastrolia and Possamaï (2018) [22], Hernández Santibáñez and Mastrolia (2019) [14], Hu, Ren, and Touzi (2019) [17], Hubert (2023) [18] or Keppo, Touzi, and Zuo (2024) [21]. Naturally, the study of applications fitting into the general framework proposed in [9] can be done in a fairly straightforward way, directly using the results, as done in [8, 1]. Nevertheless, the recent research mentioned above has also attempted to generalise principal–agent problems, especially to consider many agents (see [18, 2]), or even a continuum (see [11]). If one follows the approach developed in [9], these extensions require to introduce the so-called multidimensional or mean-field 2BSDEs, and study, or assume, the well-posedness of such systems. However, while the theory on (one-dimensional) 2BSDEs is sufficiently developed to address relatively general principal–agent problems, there is currently no definition nor existence results on multidimensional/mean-field 2BSDEs. This further motivates the presentation in this paper of an alternative approach. We believe that this novel approach, which avoids relying on the theory of 2BSDEs, will greatly facilitate the study of relevant generalised principal–agent problems, especially with many agents, or non-continuous output processes.
@@
Before introducing the approach developed in this paper, we wish to provide more intuition on the problem under consideration via a short detour through the so-called first-best case. When studying principal–agent problems with moral hazard, it is quite common to compare the results with this benchmark scenario, in which the principal can directly optimise the agent’s effort. This problem is usually easier to solve, as one only has to consider the principal’s stochastic control problem under the agent’s participation constraint. It may happen in the moral-hazard case, sometimes also referred to as second-best, that the principal can actually deduce, through her observation, the effective efforts of the agent. In this scenario, the two problems, first-best and second-best, are then equivalent if the principal can design a contract which strongly penalises the agent whenever he deviates from the optimal effort recommended by the principal. Typically, the principal will offer to the agent a minimal [^1] Minimal in the sense that it would just be sufficient to ensure the participation of agent. compensation if he follows the recommendation, and threaten him with a high penalty if he deviates from it. These forcing contracts can usually be implemented if no strong assumptions are made on the admissible contract’s form.
@@
As highlighted before, the approach developed by Cvitanić, Possamaï, and Touzi (2018) [9] relies in particular on the observation that, if the principal have access to the trajectory of the output process $X$, she can actually compute its quadratic variation $\langle X\rangle$ pathwise, and thus design contracts of the form (2.11), indexed on $X$ and $\langle X\rangle$ through sensitivity parameters $Z$ and $\Gamma$ to be optimised. The main idea behind our reformulation of the original principal–agent problem with volatility control is that, since the principal can deduce the quadratic variation $\langle X\rangle$, she could probably force the agent to choose his efforts in order to achieve an optimal quadratic variation. In other words, the idea here is to study an alternative ‘first-best’ formulation of the original problem, in which the principal enforces a contract $\xi$ as well as a quadratic variation process. The agent should then optimise his effort in order to maximise his expected utility given such contract $\xi$ as before, but now the admissible set of efforts is restricted to those achieving the quadratic variation imposed by the principal.
When considering the agent’s problem in this ‘first-best’ alternative formulation, described with more details in Section 3.1, one can remark that, since the quadratic variation is a priori fixed by the principal, the agent’s weak formulation boils down to choosing equivalent probability measures by means of Girsanov theorem. For a fixed quadratic variation process and any contract $\xi$, the agent’s continuation utility can therefore be represented by means of a BSDE, instead of a 2BSDE in the original formulation. More informally, one can actually apply the so-called ‘Sannikov’s trick’ to obtain a relevant representation of the contract $\xi$, which is then similar to the optimal form of contracts in principal–agent problems with drift control only, namely (1.1). Following this straightforward approach, we establish in Section 3.2 that the principal’s problem boils down to a more standard stochastic control problem, in the sense that she should maximise her expected utility by optimally choosing the contract’s parameters, namely a constant $y_{0}\in\mathbb{R}$ and an appropriate process $Z$, as well as the quadratic variation process.
@@
However, as mentioned before, the moral-hazard case is equivalent to its first-best counterpart if the principal can deduce the agent’s control from her observations and design a forcing contract. Indeed, even if the principal observes the agent’s control, she may not have enough bargaining power to actually force the agent to perform the optimal efforts. Similarly, in our problem, even if the principal observes the quadratic variation of the output process, she may not be able to actually force the agent to follow the recommended quadratic variation. Therefore, to ensure the equivalence between the original problem and its ‘first-best’ reformulation, we need to introduce appropriate forcing contracts, which allow the principal to achieve her ‘first-best’ utility even if she cannot directly control the quadratic variation. It is already quite clear that the contracts introduced in [9] will perfectly serve this purpose. Indeed, as these contracts are optimal when the principal observes the quadratic variation, they should allow her to achieve the expected utility she would have if she could directly control the quadratic variation process. We thus use in Section 3.3 the contract form (1.2) to conclude that the principal’s value in the reformulated problem coincides with the one derived in [9]. In other words, from the principal’s point of view, optimising over $\Gamma$ in the contract form (1.2) or directly optimising over admissible quadratic variation lead to the exact same results. In particular, this result confirms that the contracts introduced in [9] are also optimal when the agent’s efforts on the volatility can actually be determined by the principal through her observation of the quadratic variation. Interestingly, our approach further reveals the importance of an additional assumption under which the contract form (1.2) is indeed optimal and the above equivalence holds. Specifically, if Assumption B is not satisfied, the method of [9] may fail to yield optimal contracts, as explained in Remark 3.12 and illustrated through a counter-example in Section 4.3.
@@
To summarise, this paper provides an alternative yet more accessible approach to solve principal–agent problems with volatility control, in particular to prove that the contract form introduced in [9] is without loss of generality from the principal’s point of view. More precisely, we first solve a ‘first-best’ reformulation of the original problem by means of the theory of BSDEs, and then show that the contract form introduced in [9] allows to achieve this first-best scenario, and is therefore naturally optimal. We believe that this simpler approach will enable better dissemination of these problems in various fields across mathematical finance, economics, management sciences, operation research; fostering their use in many relevant applications. Furthermore, as this new approach only relies on the theory of BSDEs, we believe that it will facilitate the study of relevant extensions of principal–agent problems, for example when considering such a Stackelberg game with many agents, as well as its mean-field limit, when the principal can contract with a continuum of agents. We nevertheless emphasise that our approach relies on Assumption A, which is not assumed in [9]. More precisely, to apply the BSDE theory without introducing an orthogonal martingale component, we require a martingale representation theorem to hold; Assumption A provides a sufficient condition for this. We refer the reader to Remarks 2.2 and 3.3 for a more detailed discussion. While this assumption appears necessary for our current arguments, we believe it constitutes a relatively mild condition, especially given the significant simplification afforded by our approach. Extending the method beyond the scope of Assumption A remains an open question, but we believe there is good reason to expect that such generalisations are possible and would preserve the validity of our main results.
@@ 2 General model
In order to highlight that our alternative approach to principal–agent problems with volatility control allows to obtain the same results as in [9], we focus here on the exact same model. We believe that this model is sufficiently general to convince the reader of the relevance of our novel approach, and that it could then be extended in a (relatively) straightforward way to solve more complex problems.
@@ 3. $(iii)$
as well as for any $\nu\in\widebar{\cal U}$, the set of corresponding probability measures $\widebar{\cal P}(\nu):=\{\mathbb{P}\in\widebar{\cal P}:\nu\in\widebar{\cal U}(\mathbb{P})\}$. Finally, to avoid technical considerations (see Remark 2.2 below), we make the following standing assumption.
Assumption A.
Each agent’s admissible controls induce a unique weak solution to SDE (2.3), i.e. $\nu\in{\cal U}$ with
@@
Assumption A is not enforced in [9], but is considered here to avoid technical considerations regarding the martingale representation property. We refer to Remark 3.3 for further discussions on this point, but wish to justify here that this assumption is not far-fetched. Indeed, it is relatively common in the study of principal–agent problems to at least assume that the martingale representation is valid (see Hernández Santibáñez (2024) [13, Assumption 2.2. and Remark 2.3] for example), and weak uniqueness for SDE (2.3) is a sufficient and more explicit condition for such representation to hold. Furthermore, as one of the main objectives of principal–agent problems is to determine ‘revealing’ contracts, i.e. contracts that reveal the agent’s optimal efforts, it is actually questionable to consider efforts that would induce several weak solutions for SDE (2.3), especially in view of applications.
@@ 2.2 The original principal–agent problem
$V_{\rm A}(\xi):=\sup_{\mathbb{P}\in{\cal P}}J_{\rm A}(\xi,\mathbb{P}),\;\text{ with }\;J_{\rm A}(\xi,\mathbb{P}):=\mathbb{E}^{\mathbb{P}}\bigg{[}{\cal K}_{\rm A}^{\mathbb{P}}(T)\,\xi-\int_{0}^{T}{\cal K}_{\rm A}^{\mathbb{P}}(s)c_{\rm A}\big{(}s,X,\nu_{s}^{\mathbb{P}}\big{)}\mathrm{d}s\bigg{]},\quad\mathbb{P}\in{\cal P},$ (2.5)
@@ 1. $(i)$
    the measurable cost function $c_{\rm A}:[0,T]\times{\cal C}([0,T],\mathbb{R}^{d})\times U\longrightarrow\mathbb{R}$ is such that for any $u\in U$, the map $(t,x)\longmapsto c_{\rm A}(t,x,u)$ is $\mathbb{F}$-optional, and satisfies the following integrability condition:
    $\sup_{\mathbb{P}\in{\cal P}}\mathbb{E}^{\mathbb{P}}\bigg{[}\int_{0}^{T}\big{|}c_{\rm A}\big{(}t,X,\nu^{\mathbb{P}}_{t}\big{)}\big{|}^{p}\mathrm{d}t\bigg{]}<+\infty,\;\text{for some }p>1;$ (2.6)
@@ 2. $(ii)$
    where $k_{\rm A}:[0,T]\times{\cal C}([0,T],\mathbb{R}^{d})\times U\longrightarrow\mathbb{R}$ is assumed to be measurable, bounded, and such that for any $u\in U$, the map $(t,x)\longmapsto k_{\rm A}(t,x,u)$ is $\mathbb{F}$-optional.
@@
As mentioned in [9, Remark 2.1], the model and the approach can be extended in a straightforward way to account for risk-aversion in the agent’s utility function. More precisely, in the agent’s expected utility (2.5), $\xi$ can be replaced by $U_{\rm A}(\xi)$ for an appropriate utility function $U_{\rm A}$, modulo the modification of the integrability condition (${\rm I}^{p}_{\rm A}$). As we choose to follow the model considered in [9], we do not consider any utility functions here, but we will nevertheless informally address a case of CARA utility functions through the examples in Section 4.
@@
The principal’s optimisation problem is then defined as follows,